# GenTorrent: Scaling Large Language Model Serving with An Overlay Network

### v3

@@ 3.2 Anonymous Overlay of Users
We propose a novel approach that achieves the advantages of both Onion and sliced routing and fits the requirements of GenTorrent exactly. In GenTorrent, each user uses Onion routing to establish $n$ proxies. In this process, path failures and redundancy do not cause high resource waste because the establishment message is very short. After proxies are established, the user and model nodes rely on Garlic for prompt and response messages.
Using S-IDA. We briefly present an existing protocol Secure Information Dispersal Algorithm (S-IDA) [23] that is essential to achieve message confidentiality, anonymous routing, and failure resilience in GenTorrent. An Information Dispersal Algorithm (IDA) splits a message $M$ into multiple fragments, each of which is in ciphertext. If the receiver gets enough fragments, it can recover the plaintext message. Two most well-known IDAs are Shamir’s secret sharing (SSS) [38] and Rabin’s IDA [35]. Both of them requires $k$ out of $n$ fragments to recover the original message $M$. However, each has its weakness. SSS ensures any k1k-1 fragments provide no information about $M$, but each fragment has the same length as $M$, results in $n\times$ bandwidth cost. In Rabin’s IDA, each fragment is around $l/k$ long where $l$ is the size of $M$ but Rabin’s IDA does not guarantee that <k<k fragments provide no information about $M$. Secure IDA (S-IDA) [23, 34] improves the security level of IDA by combining it with symmetric encryption, such as Advanced Encryption Standard (AES). The sender applies the following steps.
1. Encrypts $M$ by an AES key $K$, getting $\{M\}_{K}$.
2. Splits $\{M\}_{K}$ into $n$ fragments, $M_{1},M_{2},...,M_{n}$ by a $k$-threshold Rabin’s IDA.
3. Splits $K$ into $n$ fragments, $K_{1},K_{2},...,K_{n}$ by a $k$-threshold SSS.
4. Constructs $n$ messages called cloves, $C_{1},C_{2},...,C_{n}$, where $C_{i}$ includes two fields $M_{i}$ and $K_{i}$.
5. Sends the cloves to the receiver using $n$ different paths.
By receiving at least $k$ cloves, the receiver:
1. Recovers $\{M\}_{K}$ using Rabin’s IDA and the key $K$ by SSS.
2. Decrypts $\{M\}_{K}$ and gets $M$.
In this way, only when a node collects $k$ cloves, it can recover the key and decrypted $\{M\}_{K}$. Without the complete key, it can only see the ciphertext related to $\{M\}_{K}$. We call the above process $(n,k)$ S-IDA.
Figure 2: Process of proxy establishment.
@@
We present the detailed anonymous communication protocol in GenTorrent.
Step 1. Preparation. When a user node $u$ starts running the GenTorrent client program, it joins the overlay network. The verification nodes’ IP and public keys are available to the public. The user $u$ contacts an arbitrary verification node to download a list of overlay users, called the user list, and a list of model nodes, called the mode node list. The user list is a consensus among all verification nodes and signed by more than 2/3 verification nodes. Each entry in the user list includes the public key and IP address of an overlay user. The users can be further grouped into regions, each of which has a separate list including the users around a certain network location. Users in the same region will be used to form anonymous paths. Grouping users into regions will significantly reduce the latency of anonymous paths, and a region should also include a sufficient number of users (e.g., over 1000) to ensure anonymity. In addition, GenTorrent must determine two parameters $n$ and $k$ ($n>k$) for $(n,k)$ S-IDA. It allows up to nkn-k failed paths among $n$ overlay paths.
2. Proxy and path establishment. Each user node, say $u$, needs to find $N$ proxies ($N\geq n$). To find each proxy, the user $u$ selects $l$ other users from the user list, where $l$ is the path length. Conventional wisdom indicates that anonymous paths with three hops in Tor (not including the destination) can achieve an appropriate balance between security and latency [3]. Hence GenTorrent uses $l=3$. The user $u$ then builds an Onion path using the public keys of these 3 selected users and uses the hash value of both $u$ and the last user on the path to be the path session ID $I$, as illustrated in Fig. 2. The last user $p$ on the path becomes a proxy of $u$, and every node on the path stores the predecessor and successor together with the path session ID $I$.
The above process might fail due to user dynamics but $u$ can easily try other paths and the resource loss is small, because the proxy establishment messages are short. By repeating the above process multiple times, each user $u$ can establish $N$ proxies using Onion paths. In later LLM prompts and responses, no public-key cryptographic operations are needed on the paths, because the relay users will use the stored path ID, predecessors, and successors for forwarding. In addition, different paths of $u$ has different IDs, hence colluding users on different paths cannot link the paths to the user.
3. Prompt messages. When a user node $u$ plans to send a prompt or query message $Q$ to GenTorrent, it first selects a model node from the model node list and writes its IP as the destination. As illustrated in Fig. 3, $u$ prepares $n$ slices (called cloves) of $Q$ using $(n,k)$ S-IDA and sends them to $n$ proxies respectively using the established paths. The query message $Q$ includes only the prompt and model node IP without any information about $u$. The application-layer header of each clove includes the path ID $I$ for the corresponding proxy but does not include the proxy IP. In this way, the user and proxy are still kept anonymous to the relay users. Also, the prompt messages are kept confidential. When each proxy receives the clove, it directly sends the clove to the destination model node, which is not anonymous. The prompt message also includes the IP addresses of at least $n$ proxies of $u$, which are revealed to the model nodes when it receives $\geq k$ cloves.
@@ 3.3 Overlay Forwarding among Model Nodes
All model nodes serving the same LLM form a logical group. Upon receiving a request from a user, a model node can serve this request or forward the request to other nodes in the same group. If the number of models in a group exceeds a size (50 in our implementation), the group can further be divided into multiple groups, potentially based on geographic regions.
The forwarding logic, illustrated in Fig. 5, is designed to consider both KV cache reusing and load balancing objectives. The detailed design includes the following components.
Hash-Radix tree. Recent studies [49, 41, 9] show that sharing and reusing KV cache among different GPUs can significantly reduce the average latency for LLM requests. However, prior work such as SGLang [49] and Preble [41] assumes a cloud environment that has a centralized scheduler or request router. This central entity receives all requests and distributes them to the appropriate GPUs. GenTorrent lacks such a central coordinator. The centralized scheduler [49, 41] uses a radix tree [30] for searching existing KV caches that can be reused for the current prompt, based on matching prefixes. Each tree node in a radix tree represents a prefix.
In a decentralized system like GenTorrent, naively requiring every model node to maintain a full copy of the radix tree built from the KV cache information of all model nodes in the group introduces significant scalability problems: 1) High memory overhead: each node needs to spend substantial memory space to store the aggregated radix tree, including the metadata for the KV cache from all other nodes; and 2) high communication overhead: each node needs to broadcast updates about the prefixes of its local KV cache to all other nodes in the group. To resolve these problems, we propose a new distributed data structure called a Hash-Radix tree (HR-tree), detailed below.
@@
The design of the HR-tree is motivated by cuckoo filters [16], which replace hash table entries with fingerprints to reduce memory cost. An HR-tree represents the aggregated KV cache state from all overlay model nodes in the group. A request prompt can be efficiently searched in the HR-tree. If the prompt matches a sufficiently long prefix path in the HR-tree, the HR-tree returns the model node that holds the KV cache corresponding to the matched prefix. Otherwise, the HR-tree reports no match.
@@
When each model node receives a user request that includes a prompt, it searches the prompt in the HR-tree. The search algorithm first applies pre-processing to divide the query prompt into chunks and compute their hashes. The algorithm then traverses a tree structure, starting from the root node. For each chunk of the querying prompt, the algorithm attempts to find a child node with the same hash value as the chunk. When there is no child node matching the hash value of the next chunk, the search stops. The HR-tree returns the search depth $d$. If $d$ exceeds a threshold, we consider it a match. Each tree node also includes a set of pointers, each referencing an entry of a separate table. The entry represents a model node that holds the corresponding KV cache of the prefix. Each entry in the table includes the node’s IP, current load (active requests), latency, and other information. HR-tree is distinct from the block hashing in vLLM [44, 24] for memory management and paging.
@@
Similar to Bloom and cuckoo filters [4, 16], an HR-tree might also introduce false positives: a prompt that does not match a prefix can be judged as a match by the HR-tree. For a hash value with 8 bits, the probability of a hash collision is 1/256. However, a false positive requires matching $d$ tree nodes, so the false positive rate is as small as $1/256^{d}$.
State synchronization. For each model node in a group, it periodically broadcasts the hash values of the prefixes of its local KV cache, represented as a subtree of the HR-tree, as well as other local information such as active requests that are being processed or queued, and KV cache usage. Each node aggregates the subtrees and generates an HR-tree. Using compact hashes significantly reduces the synchronization cost in this step compared to radix trees that use the entire prefixes.
Load balancing is another design consideration in addition to KV cache reusing. The main objectives of load balancing are to ensure that no model node receives more requests than its capacity, and every model node can serve a certain number of requests to establish a reputation score. Following the approach in SGLang [49], the current implementation of GenTorrent uses the number of active requests queued at a model node as the load metric. To account for the inherent hardware heterogeneity in GenTorrent, we use the relative number of requests as the metric, which is defined as the number of requests divided by the score representing a model node’s hardware capacity, ranging from 1 to 10. Moreover, GenTorrent is compatible with other metrics, such as the KV cache usage levels [41].
@@
Forwarding logic. The overlay forwarding logic, executed by every model node upon receiving a user request, determines how to forward the request, as illustrated in Fig. 5. The model node first searches the request’s prompt in its HR-tree. This search yields either a match or a miss.
If the search result is a miss, the node prioritizes load balancing. The algorithm will search its model node table, which contains the information of all model nodes in the group, collected during state synchronization. The algorithm finds the model node with the relatively fewest active requests. The user request is then forwarded to the model node.
If the search result is a match, the HR-tree identifies a list of model nodes that hold reusable KV caches. The next step is to filter this candidate list, removing any node whose current relative load is above a predefined load threshold. If there is one or more model nodes left, we pick the one with the fewest active requests and forward the current request to it. Otherwise, the logic reverts to the miss case.
@@ 3.4 Model verification
The committee of verification nodes is responsible for ensuring that the model nodes correctly run the models they promise to serve and for maintaining the reputation scores of the model nodes. The verification nodes run a BFT-like consensus protocol [7] to ensure data consistency and authenticity in the committee. The fundamental verification check is to compare a response from a target model node for a specific prompt against a reference result generated locally using the same LLM. However, verifying every user response is impractical due to scalability and potential user privacy issues.
GenTorrent introduces a sampling-based verification strategy. Instead of monitoring all responses, verification nodes periodically send challenge prompts to model nodes and verify their responses. These challenge prompts are routed by the anonymous overlay; hence, model nodes cannot distinguish verification probes from regular user requests. Verification nodes then assign each model node a reputation score (defined later) based on the quality of their response. This allows the system to evaluate the model node’s behavior probabilistically without compromising scalability or privacy.
@@
Each verification node locally deploys the same copy of the LLM as the model nodes. When receiving a challenge response from the model node, verification nodes evaluate it against their local reference result. The underlying principle is that, when using the same model and prompt, the conditional probability distribution over the next token is well-defined [15]. Suppose a verification node evaluates the correctness of a model node’s response $r$ containing $n$ tokens, $r=(t_{1},t_{2},\ldots,t_{n})$, to a challenge prompt. The verification node assesses the output token-by-token, as illustrated in Fig. 8. For each token $t_{k}$ at position $k$: 1) The verification node takes the prefix of the received response up to token $t_{k}$ and the original prompt. It uses its local LLM, conditioned on this input, to compute the probability distribution over all possible tokens at position $k+1$. 2) It then looks up the probability assigned by its local model to the actual token $t_{k+1}$ from the target model node. This probability is recorded. This process repeats for all tokens, generating a list of probabilities.
@@
Therefore, this punishment makes the reputation decrease excessively when detecting abnormal values. If a model node’s reputation drops below a critical level (like 0.4), it will be marked as untrusted.
Verification is performed epoch-by-epoch. In each epoch, the committee collaboratively assesses the correctness of model nodes and updates their reputation accordingly. GenTorrent relies on a consensus protocol Tendermint [6]. Tendermint is an open-source protocol to provide BFT consensus that ensures safety and liveness under partial synchrony. The system assumes that out of $n$ nodes, a maximum of $f$ can be Byzantine faulty, where $n\leq 3f+1$. Each update message should be signed by at least $2n/3+1$ nodes before commitment. In each epoch, a verification node is selected as the leader to send challenge requests to model nodes. To reduce overwhelming model nodes, only the leader of each epoch communicates with the model nodes, collects their responses, and broadcasts the results to the committee. Each verification node then locally verifies the results independently and votes on the updates.
@@
If the response and digital signature do not match or the model node does not respond, the leader can include an “invalid response from model node xx” message in the broadcast message, but does not reduce $x$’s reputation. It is to prevent the leader from attacking a model node by dropping or altering the response. Only when more than $1/3$ of the committee reported invalid responses for the same model node, implying that the model node misbehaves on purpose, $x$’s reputation will be reduced.
Upon receiving the information from the leader, each verification node performs an integrity check locally. It verifies whether the list of model nodes $M_{i}$ and associated prompts match the pre-determined prompt list, and checks the signatures and timestamps of each model node response. If any inconsistency is detected - e.g., different prompts or wrong signatures - verification for the current epoch is aborted. A new leader will be selected to continue verification for the next epoch. For valid challenge responses, each verification node independently computes the credibility and reputation score using its local LLM, and compares it with the score proposed by the leader $L_{i}$. The reputation update follows Tendermint’s two-phase voting protocol: Pre-Vote and Pre-Commit. If they match with negligible variance, the verification node sends a pre-vote. If the leader receives pre-votes from more than $2/3$ of the verification nodes, it proceeds to the pre-commit phase. Again, if more than $2/3$ of the nodes pre-commit, the reputation updates are finalized and broadcast to the whole system. At the end of epoch $e_{i}$, $L_{i}$ prepares the set of model nodes $M_{i+1}$ to be verified for the next epoch and the corresponding challenge prompt list.
Potential consequences for participants deploying model nodes that achieve persistently low reputation scores include suspension or revocation of their accounts in GenTorrent.
Figure 9: Anonymity vs. fraction of
malicious nodes.
Figure 10: Confidentiality vs. fraction
of malicious nodes.
Figure 11: Credit score of different
models.

### v4

@@ 3.2 Anonymous Overlay and Content Privacy
We propose a novel approach that achieves the advantages of both Onion and sliced routing and fits the requirements of PlanetServe exactly. In PlanetServe, each user uses Onion routing to establish $n$ proxies. In this process, path failures and redundancy do not cause high resource waste because the establishment message is very short. After proxies are established, the user and model nodes rely on sliced routing for prompt and response messages.
Using S-IDA. We briefly present an existing protocol Secure Information Dispersal Algorithm (S-IDA) [SIDA93] that is essential to achieve message confidentiality, anonymous routing, and failure resilience in PlanetServe. Secure IDA (S-IDA) [SIDA93, garlic] improves the security level of classic IDA [Rabin89] by combining it with symmetric encryption. The sender applies the following steps. 1) Encrypts $M$ by an AES key $K$, getting $\{M\}_{K}$. 2) Splits $\{M\}_{K}$ into $n$ fragments, $M_{1},M_{2},...,M_{n}$ by a $k$-threshold Rabin’s IDA. 3) Splits $K$ into $n$ fragments, $K_{1},...,K_{n}$ by a $k$-threshold Shamir’s secret sharing (SSS) [Shamir1979]. 4) Constructs $n$ messages called cloves, $C_{1},C_{2},...,C_{n}$, where $C_{i}$ includes two fields $M_{i}$ and $K_{i}$. 5) Sends the cloves to the receiver using $n$ different paths. By receiving at least $k$ cloves, the receiver: 1) Recovers $\{M\}_{K}$ using Rabin’s IDA and the key $K$ by SSS. 2) Decrypts $\{M\}_{K}$ and gets $M$. Only when a node collects $k$ cloves, it can recover the key and decrypted $\{M\}_{K}$. Without the complete key, it can only see the ciphertext related to $\{M\}_{K}$. We call the above process $(n,k)$ S-IDA.
@@
We present the detailed anonymous protocol.
Step 1. Preparation. A new user $u$ contacts an arbitrary verification node to download a list of overlay users, called the user list, and a list of model nodes, called the model node list, which are signed by more than 2/3 verification nodes. Each entry in the list includes the public key and IP address. PlanetServe determines $n$ and $k$ ($n>k$) for $(n,k)$ S-IDA.
2. Proxy and path establishment. Each user node, say $u$, needs to find $N$ proxies ($N\geq n$). To find each proxy, the user $u$ selects $l$ other users from the user list, where $l$ is the path length. Conventional wisdom indicates that anonymous paths with three routers in Tor can achieve an appropriate balance between security and latency [PathTor10]. Hence PlanetServe uses $l=3$. The user $u$ then builds an Onion path using the public keys of these 3 selected users and uses the hash value of both $u$ and the last user on the path to be the path session ID $I$. The last user $p$ on the path becomes a proxy of $u$, and every node on the path stores the predecessor and successor together with the path session ID $I$. The above process might fail due to user dynamics but $u$ can easily try other paths. By repeating the above process multiple times, each user $u$ can establish $N$ proxies using Onion paths. In later LLM prompts and responses, no public-key cryptographic operations are needed on the paths, because the relay users will use the stored path ID, predecessors, and successors for forwarding.
3. Prompt messages. When a user node $u$ plans to send a prompt or query message $Q$ to PlanetServe, it first selects a model node from the model node list and writes its IP as the destination. As illustrated in Fig. 2, $u$ prepares $n$ slices (called cloves) of $Q$ using $(n,k)$ S-IDA and sends them to $n$ proxies respectively using the established paths. The query message $Q$ includes only the prompt and model node IP without any information about $u$. The application-layer header of each clove includes the path ID $I$ for the corresponding proxy but does not include the proxy IP. When each proxy receives the clove, it directly sends the clove to the destination model node, which is not anonymous. The prompt message also includes the IP addresses of at least $n$ proxies of $u$, which are revealed to the model nodes when it receives $\geq k$ cloves.
@@
Content privacy. PlanetServe supports model nodes with Confidential Computing (CC) hardware, including NVIDIA Hopper and Blackwell GPUs such as H100 [NVH100, nvidia-secure-ai-blackwell-hopper]. This will allow users to choose between two privacy tiers: a default user-anonymous mode as stated above, and an optional content-privacy mode, in which both the identity and the prompt content remain protected from the node operator. For sensitive scenarios such as resume editing, healthcare, or finance, users can opt for content anonymity. Achieving content privacy protection requires hardware-based Trusted Execution Environments (TEEs) because software-only encryption can protect data in transit but not in use—once plaintext resides in GPU memory for inference, it is otherwise accessible to the host OS or hypervisor. The H100 CC mode addresses this by booting in a verified state with an on-die hardware root of trust, ensuring only NVIDIA-signed firmware is loaded, and by running inference inside an isolated GPU TEE in which all PCIe/NVLink traffic is encrypted with AES-GCM-256 and rolling IVs to prevent replay attacks. Data is exchanged with the CPU TEE (e.g., AMD SEV-SNP or Intel TDX) only through encrypted bounce buffers. Hence, plaintext is never exposed outside the enclave. Before processing, the GPU can be remotely attested via NVIDIA’s attestation service to verify its identity, firmware integrity, and CC configuration. Each LLM is launched in a Confidential VM (CVM) with the signatures of the verification committee, using the CC mode. User nodes establish end-to-end TLS sessions directly with the CVM, ensuring that neither the host nor the hypervisor can observe the prompt content. We show the performance results via implementation in Sec. 5.3.
@@ 3.3 Overlay Forwarding among Model Nodes
All model nodes serving the same LLM form a logical group. Upon receiving a request from a user, a model node can serve this request or forward the request to another node. The forwarding logic, illustrated in Fig. 4, is designed to consider both KV cache reusing and load balancing objectives. The detailed design includes the following components.
Hash-Radix tree. Recent studies [sglang24, Preble25, cheng2025distributedllm] show that sharing and reusing KV cache among different GPUs can significantly reduce the average latency for LLM requests. However, prior work such as SGLang [sglang24] and Preble [Preble25] assumes a cloud environment that has a centralized scheduler or request router. This central entity receives all requests and distributes them to the appropriate GPUs. PlanetServe lacks such a central coordinator. The centralized scheduler [sglang24, Preble25] uses a radix tree [Morrison1968] for searching existing KV caches that can be reused for the current prompt, based on matching prefixes. Each tree node in a radix tree represents a prefix.
In a decentralized system like PlanetServe, naively requiring every model node to maintain a full copy of the radix tree built from the KV cache information of all model nodes in the group introduces significant scalability problems: 1) High memory overhead: each node needs to spend substantial memory space to store the aggregated radix tree, including the metadata for the KV cache from all other nodes; and 2) high communication overhead: each node needs to broadcast updates about the prefixes of its local KV cache to all other nodes in the group. To resolve these problems, we propose a new distributed data structure called a Hash-Radix tree (HR-tree), detailed below.
@@
The design of the HR-tree is motivated by cuckoo filters [cuckoofilter2014], which replace hash table entries with fingerprints to reduce memory cost. An HR-tree represents the aggregated KV cache state from all overlay model nodes in the group. A request prompt can be efficiently searched in the HR-tree. If the prompt matches a sufficiently long prefix path in the HR-tree, the HR-tree returns the model node that holds the KV cache corresponding to the matched prefix. Otherwise, the HR-tree reports no match.
@@
When each model node receives a user request that includes a prompt, it searches the prompt in the HR-tree. The search algorithm first applies pre-processing to divide the query prompt into chunks and compute their hashes. The algorithm then traverses a tree structure, starting from the root node. For each chunk of the querying prompt, the algorithm attempts to find a child node with the same hash value as the chunk. When there is no child node matching the hash value of the next chunk, the search stops. The HR-tree returns the search depth $d$. If $d$ exceeds a threshold, we consider it a match. Each tree node also includes a set of pointers, each referencing an entry of a separate table. The entry represents a model node that holds the corresponding KV cache of the prefix. Each entry in the table includes the node’s IP, load balance factor, and reputation score.
@@
Similar to Bloom and cuckoo filters [Bloom1970, cuckoofilter2014], an HR-tree might also introduce false positives: a prompt that does not match a prefix can be judged as a match by the HR-tree. For a hash value with 8 bits, the probability of a hash collision is 1/256. However, a false positive requires matching $d$ tree nodes, so the false positive rate is as small as $1/256^{d}$.
State synchronization. For each model node in a group, it periodically broadcasts the local updates of its HR-tree, Each node keeps a recent snapshot of its HR-tree and the following updates after the snapshot. The node periodically sends a minimal but necessary update to all nodes in the domain.
Load balancing is another design consideration in addition to KV cache reusing. The main objectives of load balancing are to ensure that no model node receives more requests than its capacity, and every model node can serve a certain number of requests to establish a reputation score. In the forwarding logic shown in Fig.4, when selecting the target model nodes from a set of candidates, we always select the one with the lowest load balance (LB) factor, which is defined as $F_{LB}=L*(Q/C)$, where $L$ is the moving average of the latency to serve each request on that node, $Q$ is the number of queued requests on that node that are waiting for serving, $C$ is the capacity of the number of concurrent requests that can be served on the node. The moving average is calculated similarly to the RTT estimation with parameter $\alpha=1/8$. The LB factor is calculated locally by each node and audited by the verification committee. Each node periodically broadcasts its LB factor to other model nodes.
@@
Forwarding logic. The overlay forwarding logic, executed by every model node upon receiving a user request, determines how to forward the request, as illustrated in Fig. 4. The model node first searches the request’s prompt in its HR-tree. This search yields either a match or a miss. If the search result is a miss, the node prioritizes load balancing and forwards the request to the model node with the lowest LB factor. If the search result is a match, the HR-tree identifies a list of model nodes that hold reusable KV caches. If there is one or more model nodes left, we pick the one with the lowest LB factor and forward the current request to it.
@@ 3.4 Model verification
The committee of verification nodes is responsible for ensuring that the model nodes correctly run the models they promise to serve by maintaining their reputation scores. The verification nodes run a BFT-like consensus protocol [TendermintBFT] to ensure data consistency and authenticity in the committee. The fundamental verification check is to compare a response from a target model node for a specific prompt against a reference result generated locally using the same LLM. However, verifying every user response is impractical due to scalability and potential user privacy issues.
PlanetServe introduces a sampling-based verification strategy. Instead of monitoring all responses, verification nodes periodically send challenge prompts to model nodes and verify their responses. These challenge prompts are routed by the anonymous overlay; hence, model nodes cannot distinguish verification probes from regular user requests. Verification nodes then assign each model node a reputation score (defined later) based on the quality of their response. This allows the system to evaluate the model node’s behavior probabilistically without compromising efficiency or privacy.
@@
Each verification node locally deploys the same copy of the LLM as the model nodes. When receiving a challenge response from the model node, verification nodes evaluate it against their local reference result, which is not necessarily the same. The underlying principle is that, when using the same model and prompt, the conditional probability distribution over the next token is well-defined [fan2018hierarchical]. Suppose a verification node evaluates the correctness of a model node’s response $r$ containing $n$ tokens, $r=(t_{1},t_{2},\ldots,t_{n})$, to a challenge prompt. The verification node assesses the output token-by-token, as illustrated in Fig. 7. For each token $t_{k}$ at position $k$: 1) The verification node takes the prefix of the received response up to token $t_{k}$ and the original prompt. It uses its local LLM, conditioned on this input, to compute the probability distribution over all possible tokens at position $k+1$. 2) It then looks up the probability assigned by its local model to the actual token $t_{k+1}$ from the target model node. This process repeats for all tokens, generating a list of probabilities.
@@
Therefore, the reputation decreases excessively when detecting abnormal values. If a model node’s reputation drops below a critical level (like 0.4), it will be marked as untrusted.
Verification is performed epoch-by-epoch. In each epoch, the committee collaboratively assesses the correctness of model nodes and updates their reputation accordingly. PlanetServe relies on an open-source consensus protocol Tendermint [buchman2018latest] to ensures safety and liveness under partial synchrony. Each update message should be signed by at least $2n/3+1$ nodes before commitment. In each epoch, a verification node is selected as the leader to send challenge requests to model nodes. To reduce overwhelming model nodes, only the leader of each epoch communicates with the model nodes, collects their responses, and broadcasts the results to the committee. Each verification node then locally verifies the results independently and votes on the updates.
@@
If the response and digital signature do not match or the model node does not respond, the leader can include an “invalid response from model node xx” message in the broadcast message, but does not reduce $x$’s reputation. It is to prevent the leader from attacking a model node by dropping or altering the response. Only when more than $1/3$ of the committee reported invalid responses for the same model node, implying that $x$ misbehaves on purpose, $x$’s reputation will be reduced.
Upon receiving the information from the leader, each verification node performs an integrity check locally. It verifies whether the list of model nodes $M_{i}$ and associated prompts match the pre-determined prompt list, and checks the signatures and timestamps of each model node response. If any inconsistency is detected - e.g., different prompts or wrong signatures - verification for the current epoch is aborted. A new leader will be selected for the next epoch. For valid challenge responses, each verification node independently computes the credibility and reputation score using its local LLM, and compares it with the score proposed by the leader $L_{i}$. The reputation update follows Tendermint’s two-phase voting protocol: Pre-Vote and Pre-Commit. If they match with negligible variance, the verification node sends a pre-vote.
Figure 8: Anon. vs. malicious frac.
Figure 9: Conf. vs. malicious frac.
Figure 10: credit score over models.